A plate having an area of 0.6m2 is sliding down the inclined plane at 30° to the horizontal with a velocity of 0.36m/s. There is a cushion of fluid 1.8 mm thick between the plane and the plate. Find the viscosity of the fluid if the weight of the plate is 280N.


1
Expert's answer
2020-10-08T08:22:18-0400

To find the viscosity of the fluid, we use the formula

"F = \eta\frac{\Delta v}{\Delta x}\times A"

implying that "\eta = F \div\frac{\Delta v}{\Delta x} A"

where "F = mgsin30^{o}"

"mgsin30^{o} \div \frac{\Delta v}{\Delta x}\times A"

"= 280sin30\div \frac{0.36}{1.8\times10^{-3}}\times0.6"

"140 \div 120 = 1.17 (Pa . s)"